OpenStudy (anonymous):

show me the steps on this to have an answer of -6 11x = 7x - 24 show me the steps LOL

OpenStudy (anonymous):

subtract 7x from both sides getting 4x=-24 dividing both sides by 4 and get x=-6! :)

OpenStudy (anonymous):

11x - 7x = -24 4x = -24 x = -6

OpenStudy (anonymous):

ok got it LOL

OpenStudy (anonymous):

Check this out \begin{array}l\color{blue}{\normalsize\text{x}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{=}}\color{green}{\normalsize\text{ }}\color{blue}{\normalsize\text{-}}\color{purple}{\normalsize\text{6}}\color{purple}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

11x-7x=7x-7x-24 4x = -24 4x/4 = -24/4 x = -6.